# Victoria, BC, Canada

### v2

@@ 5.5 Learning from Other Examples
So far, we have shared that using MMR enables SE researchers to study their research topics from a more holistic perspective, expanding and complimenting methods to achieve more and novel insights. We also note from the scenarios we provide (and the paper examples we reference in Section 4.2) that mixed methods are particularly valuable in studying socio-technical aspects of SE phenomena. The research scenarios we provided help to demonstrate the broad and varied landscape of possible MMR designs in SE research. The ACM SigSoft Empirical Standards for Multiple and Mixed Methods also provides a list of example MMR papers sigsoft , but care has to be taken in using these and other examples of MMR designs as “exemplars” to guide future work. With this in mind, we next present “antipatterns” to watch out for when designing or critiquing an MMR design. An earlier version of these antipatterns also appears in the same Standards, contributed by the first author of this paper.
@@ 6 Antipatterns of Mixed Method Research Designs
The principles and research design patterns we described above can be considered as useful guidelines to follow when designing a study. And through the scenarios above, we illustrated how the principles of MMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers, authors and reviewers of papers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented, rather than inherent to the conduct of the research. We introduce presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be spotted at design time (or during review of a paper). We recommend researchers think about these antipatterns early in their research process.
@@ 6.1 Presentation Antipatterns
  • Uninvited guest or party crasher: This antipattern occurs when one of the methods used is not clearly introduced in a paper introduction or methodology and makes an unexpected entrance in the later part of the paper, such as in the discussion or limitations sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the paper, justifying why they were used. Even with emergent designs, as in the case of Sam (Sec. 5.2), all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMR when one of the approaches merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper. For example, in the case of the embedded design carried out by Zara (Sec. 5.4), if the treatment of the embedded part of the study (to find out how the tool impacted their confidence in testing) was shallow and the number of participants surveyed was very low, the paper would have oversold the contribution as a MMR study.
  • Limitation shirker: This antipattern occurs when there is a failure to discuss limitations from all methods used or from their integration in a study report. MMR will have a much longer and more in depth limitation section than a single study paper, and the limitations reported should cover all methods used and be congruent with the procedures mixed, as well as address how the data or analysis was integrated. In the case of Vicki (Sec. 5.3), failure to describe the limitations in how participants were observed (in this case, perhaps other forms of communication occurred outside Slack) would weaken the conclusions from integrating the insights from the observations with the commit behaviour.
@@ 6.2 Study Design Antipatterns
These antipatterns are ingrained in how the study was designed or conducted. They are hard to fix afterwards but thinking about them early on can help researchers avoid them.
@@
  • Missing the mark: Misalignment of MMR design with the research question or objective. This can be seen when one (or even both) of the methods do not make sense in light of the research question being asked. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked (whether planned up front or emergent). In the case of Zara, the research question considered not just the testing speed, but also student confidence. This could have been considered as two separate studies, but the second was done as an embedded study of convenience to broaden the insights collected.
@@
  • Selling your soul: This antipattern is related to “Missing the mark” and may occur when a researcher employs an additional method to appeal to a future reviewer who they believe has preference for a particular methodology, but the use of the procedure does not contribute substantively to the research findings. This can be addressed by using the method one believes in or by ensuring that the researcher draws out meta-inferences from the different procedures used. In the case of Ali (Sec. 5.1), the findings from the interviews may have been sufficient to stand on their own merit. But Ali may also have been influenced to do a survey to appease to a reviewer that may be focused on knowing if the findings generalized to more teams in the company.
@@
  • Cargo cult research [^8] Cargo Cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers. Collaborating with other researchers with broader expertise may help avoid this antipattern. For example, in the case of Vicki adding a data mining method to the observation part of the method may require working with collaborators with mining expertise if Vicki lacks that expertise.
@@
  • Sample contamination: This antipattern may occur in an MMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). For example, in the case of Ali, Ali should be careful not to survey the participants already interviewed. However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional (for example in the case of a cross-over design). Any possible contamination should be justified or reported, and mentioned in limitations.
@@
  • Lost opportunity: In an MMR sequential design, when there is a failure to use findings from an earlier study during the development of an instrument for a study in a later phase of the research. This is an issue to be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or two separate studies with two different contributions). In the case of Ali, not using insights from the interviews to ask pertinent questions in the survey about the security tool mandate could be a lost opportunity.
@@
  • Integration failure: We talked about failure to integrate data or insights as a quality attribute earlier in the paper. This antipattern occurs when there is poor integration of findings from all methods used in the study design. This can be fixed by planning how the results/findings will be integrated and making sure that the results/findings from both methods are integrated or merged in the results/findings section of the paper. In the case of Zara’s research, it may be difficult to see how to integrate the insights from the experiment and embedded survey. The insights of how the students became faster over time in the experiment, and were learning as gleaned from their self reports, can be integrated to emphasize the insight about learning using the automated testing tool and add credibility to the insights about effective testing from the dominant study. It may also be the case that the researcher integrates the findings from the mixed methods, but fails to adequately report them in the presentation (note, we could call this antipattern “Integration Presentation Failure”).
@@
  • Questionable ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve. In the cases of all the fictional researchers above, we mentioned at least some of the ethical issues to be considered. All too often though we see papers were the respondents are not treated with this same care. We refer the reader to the points raised about conducting MMR ethically in Section 3.4.
@@ 7 Discussion
In this section, we address a number of issues and nuanced related to MMR. We start by expanding on how integration in MMR can be achieved and presented. We then discuss its synergies with other methods, including difference between mixed and multi methods research. Finally, we present a call for action.
7.1 Achieving Integration
Integration serves as a foundational element of MMR, as noted in Section 2 and depicted in Figure 1, and should be thoroughly detailed in the final reports of any MMR study. Creswell and Plano Clark endorse this view (poth_sage_2023, , Chapter 2) and explain that a comprehensive description of integration, considering both its intent and procedural aspects, enhances transparency by demonstrating how different methods were effectively employed. Below we present four types of integration in MMR, as recommended by Kuckartz and Rädiker (poth_sage_2023, , Chapter 22).
  • Sequential integration occurs when the findings from one method/study inform the application of another method/study. For instance, the results of an initial quaN method may guide the development of quaL data collection instruments or inform sampling strategies for the following phase. Sequential integration is seen in Ali and Sam’s scenarios (Sections 5.1 and 5.2).
  • Results-based integration involves combining, comparing, and relating findings only after the analyses of both quaL and quaN method have been completed. This type of integration focuses on synthesizing the outcomes of each method/study (quaL and quaN) to derive comprehensive insights. Results-based integration is seen in Vicki’s scenario (Sec. 5.3).
  • Data-based integration involves the simultaneous analysis and interrelation of quaL and quaN data during the analysis phase. This type of integration requires that both data types are associated with the same underlying source, ensuring that quaL and quaN data can be distinctly assigned and connected to each individual case. This alignment allows for a more cohesive and in-depth examination of the research subject. Data-based integration is seen in Zara’s scenario (Sec. 5.4).
  • Transformation-based integration occurs when quaL data is converted into quaN data, or vice versa, allowing for integration within a single “data world” (quaL or quaN). This type of integration enables researchers to analyze transformed data using the corresponding analytical techniques, enhancing the depth and breadth of the analysis. Transformation-based integration was not covered in our scenarios but examples can be found in (poth_sage_2023, , Chapter 22) and onwuegbuzie_routledge_2021 (dedicated book on how researchers conduct analysis in MMR).
Additional examples of integration can be found in Poth’s book poth_sage_2023 , including a dedicated section on “designing innovative integrations with technology” organized by Guetterman, who introduced the idea of joint displays during integration in MMR (see guetterman_integrating_2015 ; guetterman_visuals_2021 ). Di Penta and Tamburri Dipenta also propose the adoption of mind maps to gain an overview of the design to ensure there is alignment with the study’s quaN/quaL focus, and to aid in identifying gaps for further exploration, and to address threats to validity.
@@ 7.2 Synergies with Other Methods
Mixed methods research can easily be confused with multi methods research. The difference between mixed methods and multi methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi methods can be applied with any combination including quaN-quaN or quaL-quaL. The principles we outline in Section 3 all apply to multi method research as well but we focus this paper on the use of the designs for quaN and quaL methods which can bring more emphasis on the socio-technical aspects in SE research.
@@
Finally, although the focus of this paper remains on primary studies, secondary research, such as systematic reviews of the literature (SLRs) Kitchenham2023SEGRESS ; PagePRISMA2020 , families of experiments Basili1999456 ; Santos2020566 , and rapid reviews Cartaxo2020 ; Bjarnason2023 that ground their application to methodological aspects and evidence synthesis using a mixed methods approach, can also benefit from our guidelines.
@@ 7.3 A Call to Action
In writing these guidelines, we do not simply promote MMR but emphasize using MMR ethically, as captured in the principle of ethical research. As a research community we have conducted large-scale surveys with developers and mined developer forums which in some cases has led to outrage in the developer community and introduction of new policies for unsolicited research calls for participation (see Section 3.4). As researchers, we should have empathy for the people we are studying, and indeed aim to help. They are real people. This is another reason why we include examples of fictional researchers, painting a picture of these researchers as human beings and the struggles they may face. Many of these struggles are not shared in the articles they write but perhaps could be so we can learn from each other. We recommend using MMR to make SE research more human focused – amplifying the human side of the quantitative data with the inclusion of qualitative data, both collected ethically.
As the lure of AI tools tempts the SE research community with promises of easing various research tasks and procedures, it is timely that we get into the habit of asking ourselves ‘why’, and not just be driven by the temptation of ‘why not’, as captured in our principle of methodological rationale. At the same time, as we begin to explore the risk and the opportunities of using LLMs for qualitative research bano2023LLM ; hoda2024Book , let us remind ourselves of the value of ‘deep work’, captured in the principle of procedural rigour. We strongly believe that using MMR studies conducted in interdisciplinary research settings can help address complex, socio-technical research questions and drive meaningful research, captured in our principle of novel integrated insights.
These guidelines for using MMR in SE are required now more than ever. We hope this will help present and future SE researchers use mixed methods research ethically, rationally, and rigorously to generate novel insights.
@@ 8 Conclusions
In this paper, we provide an overview of how mixed methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM Sigsoft Empirical Standards sigsoft , as well as an earlier paper by Easterbrook et al. on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm. Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the principles, characteristics and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns. There is a lot more to read and learn about mixed methods, and the references we provide too can provide a starting point for researchers to learn more.
@@
The “selecting empirical research methods” paper easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed methods paper should have a clear research question that aligns with the research design; the procedures used should answer this question; the limitations of any one method or procedure can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
@@
We also elaborate on four core principles of mixed methods research that are: providing methodological rationale for the study design; ensuring the design leads to novel integrated insights by the mixing of methods; assuring the rigorous use of individual methods and that they are mixed in a rigorous way; and finally, conducting the study in an ethical manner. These principles, at a high level, also apply to any research study (as alluded to above), but we elaborate on how these principles can be considered specifically within the context of a mixed methods research study.
@@
We further describe the core characteristics that shape a diverse landscape of mixed methods research designs. These characteristics, which include the timing of procedures, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscripts. We do not provide a comprehensive description of all possible designs, but through several research scenarios, we show how the principles and characteristics can be considered as trade-offs and design choices to be made. The integration process throughout the research design is what sets mixed methods apart.
@@
Research Paradigms
A research philosophy refers to beliefs, assumptions, values, and practices constituting a particular worldview or framework for understanding phenomena within a specific discipline or field of study (e.g., Positivism, Constructivism, Pragmatism, etc.). Paradigms provide a lens through which researchers interpret and make sense of reality, guiding their observations, theories, methods, and interpretations easterbrook_selecting_2008 , (hoda2024Book, , Chapter 5)
Strategy
Research Design
A research design involves selecting the most appropriate research design based on the nature of the research question and the desired outcomes (i.e., the researcher’s methodological choice if using the mono-method, multi-method, or mixed methods).
Data Collection Method
Researchers must decide how to gather data relevant to their research question (i.e., techniques and procedures). This may involve methods such as surveys, interviews, observations, or a combination of these approaches.
@@
Qualitative and Quantitative
@@
A multi-method study may be planned or emergent and involves either two or more quaN procedures or two or more quaL procedures, but not a mix of both quaL and quaN. The main goal is typically to expand on the insights across procedures, and they may be carried out sequentially or in parallel. Distinct research methods can be used in a multi-method design.

### v3

@@ 5.5 Learning from Other Examples
So far, we have shared that using MMR enables SE researchers to study their research topics from a more holistic perspective, expanding and complimenting methods to achieve more and novel insights. We also note from the scenarios we provide (and the paper examples we reference in Section 4.3) that mixed methods are particularly valuable in studying socio-technical aspects of SE phenomena. The research scenarios we provided help to demonstrate the broad and varied landscape of possible MMR designs in SE research. The ACM SigSoft Empirical Standards for Multiple and Mixed Methods also provides a list of example MMR papers sigsoft , but care has to be taken in using these and other examples of MMR designs as “exemplars” to guide future work. With this in mind, we next present “antipatterns” to avoid or to watch out for when designing or critiquing an MMR design. An earlier version of these antipatterns also appears in the ACM SigSoft Standards, contributed by the first author of this paper.
@@ 6 Antipatterns of Mixed Method Research Designs
The principles and research design patterns we described above can be followed when designing a study. And through the scenarios above, we illustrated how the principles of MMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers, authors and reviewers of papers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented, rather than inherent to the conduct of the research. We introduce presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be identified at design time (or during review of a paper). We recommend researchers think about these antipatterns early in their research process.
@@ 6.1 Presentation Antipatterns
  • Uninvited guest or party crasher: This antipattern occurs when one of the methods used is not clearly introduced in a paper introduction or methodology and makes an unexpected entrance in the later part of the paper, such as in the discussion or limitation sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the paper, justifying why they were used. Even with emergent designs, as in the case of Sam (Sec. 5.2), all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMR when one of the methods merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper. For example, in the case of the embedded design carried out by Zara (Sec. 5.4), if the treatment of the embedded part of the study (to find out how the tool impacted confidence in testing) was shallow and the number of participants surveyed was very low, the paper would have oversold the contribution as a MMR study.
  • Limitation shirker: This antipattern occurs when there is a failure to discuss limitations from all methods used or from their integration in a study report. MMR will have a much longer and more in depth limitation section than a single study paper, and the limitations reported should cover all methods used and be congruent with the methods mixed, as well as address how the data or analysis was integrated. In the case of Vicki (Sec. 5.3), failure to describe the limitations in how participants were observed (in this case, perhaps other forms of communication occurred outside Slack) would weaken the conclusions from integrating the insights from the observations of the commit behaviour.
@@ 6.2 Study Design Antipatterns
These antipatterns are ingrained in how the study was designed or conducted. They are hard if not impossible to fix afterwards but thinking about them early on can help researchers avoid them.
@@
  • Missing the mark: Misalignment of MMR design with the research question or objective. This can be seen when one (or even both) of the methods do not make sense in light of the research question being posed. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked. In the case of Zara, the research question considered not just the testing speed, but also testing confidence. If the embedded method did not lead to insights about testing confidence through the data integration step, this design could have been described as “missing the mark”.
@@
  • Selling your soul: This antipattern is related to “Missing the mark” and may occur when a researcher employs an additional method to appeal to a future reviewer who they believe has preference for a particular methodology or study outcome, but the use of the method does not contribute substantively to the research findings. This can be addressed by using the design one authentically believes in, or by ensuring that the researcher draws out meta-inferences from the different methods used. In the case of Ali (Sec. 5.1), the findings from the interviews may have been sufficient to stand on their own merit. But Ali may also have been influenced to do a quaNsurvey to appease to a reviewer that may be focused on knowing if the findings generalized to more teams in the company.
@@
  • Cargo cult research [^11] Cargo Cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using research methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers. Collaborating with other researchers with broader expertise may help avoid this antipattern. For example, in the case of Vicki adding a data mining method to the observation part of the design may require working with collaborators with mining expertise if Vicki lacks that expertise.
@@
  • Sample contamination: This antipattern may occur in an MMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). For example, in the case of Ali, Ali should be careful not to survey the participants already interviewed. However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional (for example, in the case of a embedded design). Any possible contamination should be justified or reported, and mentioned in limitations.
@@
  • Lost opportunity: This may occur in an MMR sequential design, when there is a failure to use findings from an earlier method during the development of an instrument for a later method. This should be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or as two separate studies with two different contributions). In the case of Ali, not using insights from the interviews to ask pertinent questions in the survey about the security tool mandate could be a lost opportunity.
@@
  • Integration failure: We talked about failure to integrate data or insights as a quality attribute earlier in the paper. This antipattern occurs when there is poor integration of findings from all methods used in the study design. This can be fixed by planning how the results/findings will be integrated and making sure that the results/findings from both methods are integrated or merged in the results/findings section of the paper (or earlier in the case of an embedded design). In the case of Zara’s research, it may be difficult to see how to integrate the insights from the experiment and embedded survey. The insights of how the students became faster over time in the experiment, and were learning as gleaned from their self reports, can be integrated to emphasize the insight about learning using the automated testing tool and add credibility to the insights about effective testing from the dominant study. It may also be the case that the researcher integrates the findings from the mixed methods, but fails to adequately report them in the presentation (note, we could call this antipattern “Integration Presentation Failure”).
@@
  • Questionable ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve. In the cases of all the fictional researchers above, we mentioned at least some of the ethical issues to be considered. All too often though we see papers were the respondents are not treated with this same care. We refer the reader to the points raised about conducting MMR ethically in Section 3.4 and encourage all studies should undergo an ethical review.
@@ 7 Discussion
In this section, we address a number of issues and nuances related to MMR. We discuss MMR synergies with other methods, including difference between mixed and multi methods research. Finally, we present a call for action.
@@ 7.1 Synergies with Other Methods
Mixed methods research can easily be confused with multi methods research [^12] We remind the reader this and other terms are defined in the Glossary found in the Appendix. The difference between mixed methods and multi methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi methods can be applied with any combination of quaN-quaN or quaL-quaL methods. The principles we outline in Section 3 all apply to multi method research as well but we focus this paper on the use of the designs for quaN and quaL methods which can bring more emphasis on the socio-technical aspects in SE research.
@@
Finally, although the focus of this paper remains on primary studies, secondary research, such as systematic reviews of the literature (SLRs) Kitchenham2023SEGRESS ; PagePRISMA2020 , families of experiments Basili1999456 ; Santos2020566 , and rapid reviews Cartaxo2020 ; Bjarnason2023 that ground their application to methodological aspects and evidence synthesis using a mixed methods approach, can also benefit from the principles we synthesize in this paper.
@@ 7.2 A Call to Action
Much of SE research is inherently socio-technical and as such can benefit from the use of Mixed Methods to bring insights on the underlying context and to strengthen the theoretical contributions from our work, and capability to generalize or translate our findings to other contexts. We call for a principled use of mixed methods in SE research. Figure 7 summarizes the key points we raised in our paper: the MMR definition, a landscape of MMR designs, four principles to guide MMR in SE, and a list of antipatterns of MMR designs to be considered.
Figure 7: Mixed methods research in SE: definition, design landscape, principles and antipatterns.
In adapting the guiding principles we propose for an SE context, we do not simply promote MMR but emphasize using MMR ethically, as captured in the principle of ethical research. As a research community we have conducted large-scale surveys with developers and mined developer forums which in some cases has led to outrage in the developer community and introduction of new policies for unsolicited research calls for participation (see Section 3.4). As researchers, we should have empathy for the people we are studying, and indeed aim to help. They are real people. This is another reason why we include examples of fictional researchers, painting a picture of these researchers as human beings and the struggles they may face. Many of these struggles are not shared in the articles they write but perhaps could be so we can learn from each other. We recommend using MMR to make SE research more human focused–amplifying the human side of the quantitative data with the inclusion of qualitative data, both collected ethically.
As the lure of AI tools tempts the SE research community with promises of easing various research tasks and methods, it is timely that we get into the habit of asking ourselves ‘why’, and not just be driven by the temptation of ‘why not’, as captured in our principle of methodological rationale. At the same time, as we begin to explore the risk and the opportunities of using LLMs for qualitative research bano2023LLM ; hoda2024Book , let us remind ourselves of the value of ‘deep work’, captured in the principle of procedural rigour. We strongly believe that using MMR studies conducted in interdisciplinary research settings can help address complex, socio-technical research questions and drive meaningful research, captured in our principle of novel integrated insights.
These guiding principles for using MMR in SE are required now more than ever. We hope this first step will help present and future SE researchers use mixed methods research ethically, rationally, and rigorously to generate novel insights, and encourage our community to develop more prescriptive guidelines and playbooks for using and evaluating mixed methods research in SE.
@@ 8 Conclusions
In this paper, we provide an overview of how mixed methods research designs can be used in software engineering, with the intention to inspire their use. Other resources, such as the ACM Sigsoft Empirical Standards sigsoft , as well as an earlier paper by Easterbrook et al. on selecting a research method in software engineering easterbrook_selecting_2008 , only lightly touch on the concepts and challenges of using mixed methods research in software engineering. As can be seen from Section 2, there is a long history of research behind the mixed method paradigm and use in other disciplines, some related to SE venkatesh2013bridging . Deciding what to include (and exclude) in this paper from this large body of research was challenging and we focused on the guiding principles, properties and designs that we hope will support software engineering researchers in applying mixed methods to lead to impactful insights. Still, applying mixed methods is not a trivial exercise, as we try to demonstrate through the identification of many possible antipatterns. There is a lot more to read and learn about mixed methods, and the references we provide too can provide a starting point for researchers to learn more.
@@
The “selecting empirical research methods” paper easterbrook_selecting_2008 notably elaborates on topics such as theoretical stance and the role of theory in quantitative, qualitative and mixed research. The main takeaways from that paper also apply here: a mixed methods paper should have a clear research question that aligns with the research design; the methods used should answer this question; the limitations of any one method can be used to motivate the choice of other methods, but the limitations of and threats associated with mixing should also be reported; and the role of theory and whether the research is inductive or deductive should drive the research design and shape the integrated findings reported.
@@
We also elaborate on four core guiding principles of mixed methods research that are: providing methodological rationale for the study design; ensuring the design leads to novel integrated insights by the mixing of methods; assuring the rigorous use of individual methods and that they are mixed in a rigorous way; and finally, conducting the study in an ethical manner. These principles, at a high level, also apply to any research study (as alluded to above), but we elaborate on how these principles can be considered specifically within the context of a mixed methods research study in SE.
@@
We further describe the core properties that shape a diverse landscape of mixed methods research designs. These properties, which include the timing of methods, the inductive or deductive nature of the research, the research design planning or emergence, how data is integrated, and the mixed design, should be carefully considered up front and throughout the research process, and presented clearly in any research manuscripts. We do not provide a comprehensive description of all possible designs, but through several research scenarios, we show how the principles and properties can be considered as trade-offs and design choices to be made. The integration process throughout the research design is what sets mixed methods apart.
@@
Data Collection Method
Researchers must decide how to gather data relevant to their research question (i.e., techniques and methods). This may involve methods such as surveys, interviews, observations, or a combination of these approaches.
@@
A multi-method study may be planned or emergent and involves either two or more quaN methods or two or more quaL methods, but not a mix of both quaL and quaN. The main goal is typically to expand on the insights across methods, and they may be carried out sequentially or in parallel. Distinct research methods can be used in a multi-method design.
Qualitative and Quantitative
Research Design
A research design involves selecting the most appropriate research design based on the nature of the research question and the desired outcomes (i.e., the researcher’s methodological choice if using mono-method, multi-method, or mixed methods).
@@
Research Paradigms
A research paradigm is a framework of beliefs and practices that defines a worldview for understanding phenomena within a discipline, shaping research through ontological, epistemological, and methodological perspectives. Quantitative research, rooted in positivism, seeks objective, measurable realities; qualitative research, aligned with constructivism, explores subjective, socially constructed realities; and mixed methods research, guided by pragmatism, integrates both to address complex questions. Together, these form a trilogy of major research paradigms Johnson2004 , Kuhn1977 .
Research Strategy